OpenStudy (anonymous):

divide and simply: (4a)/(7b^2)÷(5a^2)/(8b) a. 32/35ab b. 2/35ab c. 5a^3/14b^3 d. 32ab/35

OpenStudy (anonymous):

\[\frac{4a}{7b ^{2}} \div \frac{5a ^{2}}{8b}\] \[= \frac{4a}{7b^{2}} \times \frac{8b}{5a^{2}}\] \[= \frac{4}{7b} \times \frac{8}{5a}\] \[=\frac{32}{35ab}\]

OpenStudy (anonymous):

thank you soo much @tiaph :)